# Estimating the Conformal Prediction Threshold from Noisy Labels

### v1

@@ 3.4 A More General Noise Model
Next, we will extend our approach to a more general noise model. We will assume that the noisy label $\tilde{y}$ is independent of $x$ given $y$. We also assume that the noise matrix $P(i,j)=p(\tilde{y}=j|y=i)$ is known and that the matrix P is invertible. For each $q$ define the following matrices for the clear and the noisy data: ${M}_{q}^{c}(\ell,i)=p(\ell\in C_{q}(x),y=i)$ and ${M}_{q}(\ell,i)=p(\ell\in C_{q}(x),\tilde{y}=i)$. Assuming that, given the true label $y$, the r.v. $x$ and $\tilde{y}$ are independent, we obtain:
@@
$F^{c}(q)=p(y\in C_{q}(x))=p(y\in C_{q}(x))=\sum_{i}p(i\in C_{q}(x),y=i)=\sum_{i}M^{c}_{q}(i,i)=\textrm{Tr}(M_{q}P^{-1}).$ (16)
We can estimate matrix $M_{q}$ from the noisy samples:
@@
Thus in the case of a uniform noise the coverage estimation (18) coincides with (8). If the noise matrix is unknown, it can be estimated from the noisy-label data during training [24, 9, 10]. The NACP method for a noise matrix model is summarized in Algorithm 2.
We can extend the finite sample term $\Delta$ that was developed for a uniform noise to obtain a theoretical coverage guarantee for a noise matrix model (see Section A.2). However, this approach yields large prediction sets especially in tasks with many classes and thus is ineffective. In the experiment section we show that in practice, even without adding finite sample terms, we obtain the required coverage probability.
@@ 4 Experiments
In this section, we evaluate the capabilities of our NACP algorithm on various medical and scenery imaging datasets.
@@
Compared methods. Our method takes an existing conformity score $S$ and computes a threshold $q$ that takes into account the label noise level. We implemented three popular conformal prediction scores, namely APS [16], RAPS [2] and HPS [19]. For each score $S$, we compared the following four CP methods: (1) CP (Oracle) - using a validation set with clean labels, (2) Noisy-CP - applying a standard CP on noisy labels without any modifications [3], (3) NR-CP [15] applying standard CP on an estimation of the noise-free score (4) Adaptive Conformal Classification with Noisy labels (ACNL) [17] and (5) NACP - our approach (6) NACP (w/o $\Delta$) - our approach without the finite sample coverage guarantee $\Delta$. For methods (3) and (4) we used their official codes  [^1] https://github.com/cobypenso/Noise-Robust-Conformal-Prediction 222 https://github.com/msesia/conformal-label-noise. We share our code for reproducibility [^3] https://github.com/cobypenso/Noise-Aware-Conformal-Prediction.
@@
Table 1: APS and RAPS (randomized versions) calibration results for $1\!-\!\alpha$ = 0.9 and $\epsilon=0.2$. We report the mean and the std over 1000 different splits. Bold for best result with theoretical guarantees.
Table 1 rand-APS rand-RAPS
Dataset CP Method size $\downarrow$ coverage (%) size $\downarrow$ coverage (%)
TissueMNIST CP (Oracle) 2.33 $\pm$ 0.01 90.01 $\pm$ 0.23 2.09 $\pm$ 0.01 90.01 $\pm$ 0.23
Noisy-CP 4.33 $\pm$ 0.04 98.98 $\pm$ 0.07 4.29 $\pm$ 0.04 99.12 $\pm$ 0.06
NR-CP 3.19 $\pm$ 0.02 96.09 $\pm$ 0.09 3.22 $\pm$ 0.02 96.98 $\pm$ 0.08
ACNL 3.07 $\pm$ 0.04 95.55 $\pm$ 0.24 2.52 $\pm$ 0.04 91.84 $\pm$ 0.42
NACP 2.50 $\pm$ 0.02 91.67 $\pm$ 0.27 2.42 $\pm$ 0.01 92.96 $\pm$ 0.08
NACP (w/o $\Delta$)   2.34 $\pm$ 0.02   90.05 $\pm$ 0.27   2.32 $\pm$ 0.01   92.18 $\pm$ 0.09
OrganSMNIST CP (Oracle) 1.63 $\pm$ 0.02 90.0 $\pm$ 0.58 1.63 $\pm$ 0.02 90.02 $\pm$ 0.58
Noisy-CP 5.62 $\pm$ 0.17 99.70 $\pm$ 0.06 5.75 $\pm$ 0.16 99.43 $\pm$ 0.09
NR-CP 2.88 $\pm$ 0.06 97.95 $\pm$ 0.14 3.82 $\pm$ 0.13 98.98 $\pm$ 0.12
ACNL 2.91 $\pm$ 0.26 97.92 $\pm$ 0.53 2.90 $\pm$ 0.28 97.91 $\pm$ 0.53
NACP 1.95 $\pm$ 0.06 94.03 $\pm$ 0.61 1.95 $\pm$ 0.05 94.04 $\pm$ 0.59
NACP (w/o $\Delta$)   1.63 $\pm$ 0.03   90.15 $\pm$ 0.70   1.63 $\pm$ 0.03   90.02 $\pm$ 0.67
OrganCMNIST CP (Oracle) 1.75 $\pm$ 0.04 90.04 $\pm$ 0.60 1.19 $\pm$ 0.01 90.06 $\pm$ 0.58
Noisy-CP 6.07 $\pm$ 0.19 99.96 $\pm$ 0.02 5.54 $\pm$ 0.19 99.80 $\pm$ 0.04
NR-CP 2.39 $\pm$ 0.03 98.39 $\pm$ 0.15 3.26 $\pm$ 0.99 99.70 $\pm$ 0.08
ACNL 2.23 $\pm$ 0.09 97.56 $\pm$ 0.56 1.62 $\pm$ 0.07 97.56 $\pm$ 0.55
NACP 1.92 $\pm$ 0.04 93.99 $\pm$ 0.60 1.33 $\pm$ 0.02 93.97 $\pm$ 0.60
NACP (w/o $\Delta$) 1.74 $\pm$ 0.04 90.15 $\pm$ 0.67 1.18 $\pm$ 0.02 90.12 $\pm$ 0.69
OrganAMNIST CP (Oracle) 1.03 $\pm$ 0.01 90.04 $\pm$ 0.38 1.02 $\pm$ 0.01 90.02 $\pm$ 0.37
Noisy-CP 5.45 $\pm$ 0.12 100.0 $\pm$ 0.00 5.45 $\pm$ 0.12 99.95 $\pm$ 0.01
NR-CP 1.40 $\pm$ 0.01 98.76 $\pm$ 0.09 4.54 $\pm$ 1.02 99.94 $\pm$ 0.02
ACNL 1.17 $\pm$ 0.01 95.53 $\pm$ 0.38 1.15 $\pm$ 0.01 95.53 $\pm$ 0.39
NACP 1.08 $\pm$ 0.01 92.56 $\pm$ 0.40 1.07 $\pm$ 0.01 92.55 $\pm$ 0.40
NACP (w/o $\Delta$) 1.03 $\pm$ 0.01 90.22 $\pm$ 0.44 1.02 $\pm$ 0.01 90.20 $\pm$ 0.44
PathMNIST CP (Oracle) 1.10 $\pm$ 0.02 90.03 $\pm$ 0.47 1.03 $\pm$ 0.01 90.0 $\pm$ 0.45
Noisy-CP 4.69 $\pm$ 0.13 100.00 $\pm$ 0.00 4.55 $\pm$ 0.12 99.95 $\pm$ 0.02
NR-CP 1.43 $\pm$ 0.01 98.24 $\pm$ 0.12 1.53 $\pm$ 0.03 99.00 $\pm$ 0.09
ACNL 1.22 $\pm$ 0.02 94.95 $\pm$ 0.49 1.30 $\pm$ 0.04 95.35 $\pm$ 0.55
NACP 1.18 $\pm$ 0.02 93.53 $\pm$ 0.48 1.11 $\pm$ 0.01 93.51 $\pm$ 0.47
NACP (w/o $\Delta$)   1.10 $\pm$ 0.02   90.25 $\pm$ 0.52   1.03 $\pm$ 0.01   90.24 $\pm$ 0.53
Table 2: Finite sample correction terms $\Delta$ of NACP and ACNL [17] for several medical imaging datasets and two noise levels, $n$ is the size of the validation set.
NACP ACNL
Dataset $n$classes $\epsilon=0.1$ $\epsilon=0.2$ $\epsilon=0.1$ $\epsilon=0.2$
TissueMNIST 35460 8 0.0132 0.0162 0.0158 0.0294
OrganSMNIST 5640 11 0.0331 0.0406 0.0378 0.0721
OrganCMNIST 5304 11 0.0341 0.0419 0.0380 0.0721
OrganAMNIST 12135 11 0.0225 0.0277 0.0263 0.0509
PathMNIST 8592 9 0.0268 0.0329 0.0426 0.0824
@@
$\textrm{size}=\frac{1}{n}\sum_{i}|C(x_{i})|,\hskip 28.45274pt\textrm{coverage}=\frac{1}{n}\sum_{i}{\bf 1}(y_{i}\in C(x_{i}))$
@@
Datasets. We present results on several publicly available medical and scenary imaging and classification datasets. TissuMNIST [22, 23]: This dataset contains 236,386 human kidney cortex cells, organized into 8 categories. Each gray-scale image is $32\times 32\times 7$ pixels. The 2D projections were obtained by taking the maximum pixel value along the axial-axis of each pixel, and were resized into $28\times 28$ gray-scale images [20]. PathMNIST [23]: This dataset contains 97,176 images of colon pathology with nine classes. The image size is $28\times 28$. Here, we used a train/validation/test split of 89,996/3,590/3,590 images. OrganSMNIST [23]: This dataset contains 25,221 images of abdominal CT in eleven classes. The images are $28\times 28$ in size. Here, we used a train/validation/test split of 13,940/2,452/8,829 images. OrganAMNIST and OrganCMNIST are simlar datasets, the differences of Organ{A,C,S}MNIST are the views and dataset size. We also show results on four standard scenery image datasets CIFAR-10, CIFAR-100 [8], Tiny-ImageNet, and ImageNet.
@@
Implementation details. Each task was trained by fine-tuning on a pre-trained ResNet-18 [7] network. The models were taken from the PyTorch site444 https://pytorch.org/vision/stable/models.html. We selected this network architecture because of its widespread use in classification problems. The last fully connected layer output size was modified to fit the corresponding number of classes for each dataset. For PathMNIST, TissueMNIST, OrganSMNIST, OrganCMNIST, and OrganAMNIST we used publicly available checkpoints [^5] https://github.com/MedMNIST/MedMNIST. Also for the standard dataset evaluated in Table 3 we used publicly available checkpoints. For each dataset, we combined the validation and test sets and then constructed 1000 different splits where 50% was used for the calibration phase and 50% was used for testing. In all our experiments we used $\delta=0.001$.
Table 3: CP calibration results on CIFAR-10, CIFAR-100, Tiny-ImageNet and ImageNet for $1\!-\!\alpha$ = 0.9 and $\epsilon=0.2$. We report the mean and the std over 1000 different splits. Bold for best result with theoretical guarantees.
Table 3 CIFAR-10 (10 classes) CIFAR-100 (100 classes) Tiny-ImageNet (200 classes) ImageNet (1000 classes)
rand-APS CP (Oracle) 1.1 $\pm$ 0.01 90.0 $\pm$ 0.62 6.5 $\pm$ 0.20 90.0 $\pm$ 0.43 14.9 $\pm$ 0.60 90.0 $\pm$ 0.61 16.6 $\pm$ 0.33 90.0 $\pm$ 0.26
Noisy-CP 5.1 $\pm$ 0.18 99.9 $\pm$ 0.04 50.5 $\pm$ 1.29 99.8 $\pm$ 0.04 99.7 $\pm$ 3.67 99.7 $\pm$ 0.08 502.6 $\pm$ 8.56 99.9 $\pm$ 0.01
NR-CP 2.2 $\pm$ 0.07 98.9 $\pm$ 0.12 37.6 $\pm$ 0.86 99.6 $\pm$ 0.06 79.6 $\pm$ 2.82 99.3 $\pm$ 0.11 275.6 $\pm$ 27.1 99.7 $\pm$ 0.06
ACNL 1.5 $\pm$ 0.06 96.0 $\pm$ 0.61 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00
NACP 1.3 $\pm$ 0.04 94.4 $\pm$ 0.62 9.0 $\pm$ 0.46 93.0 $\pm$ 0.49 22.6 $\pm$ 1.87 93.7 $\pm$ 0.71 20.9 $\pm$ 0.72 91.9 $\pm$ 0.32
NACP (w/o $\Delta$)   1.1 $\pm$ 0.02   90.1 $\pm$ 0.70   6.4 $\pm$ 0.28   89.9 $\pm$ 0.54   14.0 $\pm$ 0.91   89.4 $\pm$ 0.81   16.7 $\pm$ 0.51   90.0 $\pm$ 0.34
rand-RAPS CP (Oracle) 1.1 $\pm$ 0.01 90.0 $\pm$ 0.61 4.0 $\pm$ 0.08 90.0 $\pm$ 0.43 6.9 $\pm$ 0.19 90.0 $\pm$ 0.62 6.3 $\pm$ 0.06 90.0 $\pm$ 0.27
Noisy-CP 5.1 $\pm$ 0.18 99.9 $\pm$ 0.04 50.5 $\pm$ 1.33 99.8 $\pm$ 0.03 101.4 $\pm$ 3.58 99.5 $\pm$ 0.09 501.6 $\pm$ 8.51 99.9 $\pm$ 0.01
NR-CP 3.3 $\pm$ 0.20 99.6 $\pm$ 0.08 50.1 $\pm$ 1.08 99.8 $\pm$ 0.03 100.6 $\pm$ 2.88 99.5 $\pm$ 0.09 455.2 $\pm$ 20.7 99.9 $\pm$ 0.02
ACNL 1.3 $\pm$ 0.03 94.6 $\pm$ 0.65 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00
NACP 1.3 $\pm$ 0.04 94.5 $\pm$ 0.62 4.8 $\pm$ 0.13 93.0 $\pm$ 0.48 9.0 $\pm$ 0.50 93.6 $\pm$ 0.70 7.1 $\pm$ 0.13 91.9 $\pm$ 0.34
NACP (w/o $\Delta$)   1.1 $\pm$ 0.02   90.1 $\pm$ 0.69   4.0 $\pm$ 0.11   89.9 $\pm$ 0.55   6.7 $\pm$ 0.27   89.3 $\pm$ 0.80   6.3 $\pm$ 0.10   90.0 $\pm$ 0.36
HPS CP (Oracle) 0.9 $\pm$ 0.01 90.0 $\pm$ 0.59 2.0 $\pm$ 0.03 90.0 $\pm$ 0.43 3.8 $\pm$ 0.13 90.02 $\pm$ 0.58 3.6 $\pm$ 0.07 90.0 $\pm$ 0.28
Noisy-CP 5.1 $\pm$ 0.18 99.8 $\pm$ 0.04 50.1 $\pm$ 1.34 99.9 $\pm$ 0.02 98.3 $\pm$ 3.80 99.8 $\pm$ 0.05 501.3 $\pm$ 10.2 100.0 $\pm$ 0.01
NR-CP 1.2 $\pm$ 0.01 97.8 $\pm$ 0.12 11.7 $\pm$ 0.37 98.9 $\pm$ 0.08 28.1 $\pm$ 1.20 98.1 $\pm$ 0.15 55.9 $\pm$ 0.87 99.1 $\pm$ 0.02
ACNL 1.1 $\pm$ 0.03 96.0 $\pm$ 0.59 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00
NACP 1.1 $\pm$ 0.01 95.9 $\pm$ 0.18 2.5 $\pm$ 0.09 93.0 $\pm$ 0.52 7.0 $\pm$ 0.87 93.6 $\pm$ 0.72 4.8 $\pm$ 0.23 91.9 $\pm$ 0.36
NACP (w/o $\Delta$)   0.9 $\pm$ 0.02   90.0 $\pm$ 0.75   2.0 $\pm$ 0.06   89.9 $\pm$ 0.56   3.5 $\pm$ 0.24   89.3 $\pm$ 0.80   3.6 $\pm$ 0.14   90.0 $\pm$ 0.38

### v2

@@ 3.4 A More General Noise Model
The focus of this paper is on the case of a uniform noise. Our approach, however, is easily extended to a more general noise model. We will assume that the noisy label $\tilde{y}$ is independent of $x$ given $y$. We also assume that the noise matrix $P(i,j)=p(\tilde{y}=j|y=i)$ is known and that the matrix P is invertible. For each $q$ define the following matrices for the clear and the noisy data: ${M}_{q}^{c}(\ell,i)=p(\ell\in C_{q}(x),y=i)$ and ${M}_{q}(\ell,i)=p(\ell\in C_{q}(x),\tilde{y}=i)$. Assuming that, given the true label $y$, the r.v. $x$ and $\tilde{y}$ are independent, we obtain:
@@
$F^{c}(q)=p(y\in C_{q}(x))=\sum_{i}p(i\in C_{q}(x),y=i)=\sum_{i}M^{c}_{q}(i,i)=\textrm{Tr}(M_{q}P^{-1}).$ (14)
We can estimate the matrix $M_{q}$ from the noisy samples:
@@
Thus in the case of a uniform noise the coverage estimation (16) coincides with (8).
4 Related Work
In this section we review two closely related works that address the same problem of calibration with noisy labels (Sesia et al., 2024; Clarkson et al., 2024). The derivation of the noisy conformal threshold in these two works is similar to ours. These two methods compute the same threshold $q$ that satisfies $\hat{f}^{c}(q)=1-\alpha$ (16). The only minor difference is that in these two studies they use the distribution of correct labels given the noisy labels, while we use the more natural distribution of the noisy labels given the correct label. As a result, they need to know the marginal class frequencies for both the clean and noisy labels, whereas we do not. Each one of the two methods provides a different finite coverage guarantee in the form of:
$p(y\in C_{q}(x))\geq 1-\alpha-\Delta$
where $\Delta$ depends on the calibration set size $n$, the number of classes $k$, and the noise model, but it doesn’t depend on the validation dataset itself.
[…]
$\Delta$ $=c(n)+\frac{2\max_{i\in[k]}\sum_{l\neq i}|V_{il}|+\sum_{i=1}^{k}|\rho_{i}-\tilde{\rho}_{i}|}{\sqrt{n_{*}}}$ (17)
$\cdot\min\left(k^{2}\sqrt{\frac{\pi}{2}},\frac{1}{\sqrt{n^{*}}}+\sqrt{\frac{\log(2k^{2})+\log(n^{*})}{2}}\right).$
It can be easily verified that $\Delta=O(\log{k})$ and therefore the bound becomes less effective for large values of $k$.
The finite sample term $\Delta$ derived in (Clarkson et al., 2024) is:
$\Delta=\sum_{i=1}^{k}(|w_{i}^{(1)}|b(n,i)+\sum_{i\neq j}|w_{ij}^{(2)}|b(n,j))$ (18)
where $k$ is the number of classes, $w^{(1)}_{i}=P_{i,i}^{-1}\rho_{i}-\tilde{\rho}_{i}$, $w^{(2)}_{ij}=\rho_{i}P^{-1}_{ji}$, and $b(n,j)=(1-\tilde{\rho}_{j})^{n}+\sqrt{\frac{\pi}{n\tilde{\rho}_{j}}}$. $\rho_{i}=p(y=i)$ and $\tilde{\rho}_{i}=p(\tilde{y}=i)$ are the marginal clean and contaminated label probabilities and $P_{ji}=p(y=j|\tilde{y}=i)$ is the conditional label noise distribution. It can be easily verified that $\Delta=O(\sqrt{k})$ and therefore the bound becomes less effective for large values of $k$.
We note that our finite sample term $\Delta$ (see Lemma 4) does not depend on the number of classes $k$. Therefore, unlike the algorithms of (Sesia et al., 2024) and (Clarkson et al., 2024), it remains effective even in tasks with many classes. A further distinction between our approach and previous works is that their finite sample coverage guarantee is established based on the average of all the noisy calibration sets. In contrast, our approach provides an individual coverage guarantee for nearly all ($1-\delta$) of the sampled noisy calibration sets. In Section 5 we show that the average coverage guarantee obtained by (Sesia et al., 2024) and (Clarkson et al., 2024) implies that in tasks with a large number of classes, the prediction set should include all the classes and therefore it is useless. In contrast, our individual finite set coverage guarantee, on to $(1-\delta)$ portion of the noisy calibration sets, remains effective for tasks with many classes.
We note that our observation that the finite sample correction term does not depend on the number of classes, applies to the case of a uniform label noise. In the case of a general noise matrix, all finite sample correction terms are not effective.
@@ 5 Experiments
In this section, we evaluate the capabilities of our NACP algorithm on various publicly available image classification datasets.
@@
Compared methods. We implemented three popular conformal prediction scores, namely APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2021) and HPS (Vovk et al., 2005). For each score, we implemented the following CP methods: (1) CP (oracle) - using a calibration set with clean labels, (2) Noisy-CP - applying a standard CP on noisy labels without any modifications (Einbinder et al., 2022), (3) NR-CP (w/o $\Delta$) - Noise-Robust CP approach without the finite sample coverage guarantee $\Delta$, see Eq. (16) and (Sesia et al., 2024; Clarkson et al., 2024). We also implemented three methods that add finite sample coverage guarantee terms to the NR-CP method. (4) Adaptive Conformal Classification with Noisy labels (ACNL) (Sesia et al., 2024), (5) Contamination Robust Conformal Prediction (CRCP) (Clarkson et al., 2024), and (6) NACP - our approach. For methods (4), (5), we used their official codes 111 https://github.com/msesia/conformal-label-noise 222 https://github.com/jase-clarkson/cp_under_data_contamination and we share our code for reproducibility [^3] https://github.com/cobypenso/Noise-Aware-Conformal-Prediction.
@@
Table 1: CP calibration results for $1\!-\!\alpha$ = 0.9 and noise level $\epsilon=0.2$. We report the mean and the std over 1000 different splits. We show the best result with theoretical guarantees in bold.
APS RAPS HPS
Dataset CP Method size $\downarrow$ coverage(%) size $\downarrow$ coverage(%) size $\downarrow$ coverage(%)
CP (oracle) 1.1 $\pm$ 0.01 90.0 $\pm$ 0.62 1.1 $\pm$ 0.01 90.0 $\pm$ 0.61 0.9 $\pm$ 0.01 90.0 $\pm$ 0.59
Noisy-CP 5.1 $\pm$ 0.18 99.9 $\pm$ 0.04 5.1 $\pm$ 0.18 99.9 $\pm$ 0.04 5.1 $\pm$ 0.18 99.8 $\pm$ 0.04
NR-CP (w/o $\Delta$)   1.1 $\pm$ 0.02   90.1 $\pm$ 0.70   1.1 $\pm$ 0.02   90.1 $\pm$ 0.69   0.9 $\pm$ 0.02   90.0 $\pm$ 0.75
CIFAR-10 ACNL 1.5 $\pm$ 0.06 96.0 $\pm$ 0.61 1.3 $\pm$ 0.03 94.6 $\pm$ 0.65 1.1 $\pm$ 0.03 96.0 $\pm$ 0.59
(10 classes) CRCP 1.2 $\pm$ 0.03 93.7 $\pm$ 0.62 1.2 $\pm$ 0.03 93.7 $\pm$ 0.62 1.1 $\pm$ 0.01 95.7 $\pm$ 0.18
NACP 1.3 $\pm$ 0.04 94.4 $\pm$ 0.62 1.3 $\pm$ 0.04 94.5 $\pm$ 0.62 1.1 $\pm$ 0.01 95.9 $\pm$ 0.18
CP (oracle) 6.5 $\pm$ 0.20 90.0 $\pm$ 0.43 4.0 $\pm$ 0.08 90.0 $\pm$ 0.43 2.0 $\pm$ 0.03 90.0 $\pm$ 0.43
Noisy-CP 50.5 $\pm$ 1.29 99.8 $\pm$ 0.04 50.5 $\pm$ 1.33 99.8 $\pm$ 0.03 50.1 $\pm$ 1.34 99.9 $\pm$ 0.02
NR-CP (w/o $\Delta$) 6.4 $\pm$ 0.28   89.9 $\pm$ 0.54 4.0 $\pm$ 0.11   89.9 $\pm$ 0.55   2.0 $\pm$ 0.06   89.9 $\pm$ 0.56
CIFAR-100 ACNL 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00 100.0 $\pm$ 0.00
(100 classes) CRCP 25.7 $\pm$ 3.71 98.7 $\pm$ 0.39 8.5 $\pm$ 0.41 98.3 $\pm$ 0.16 11.1 $\pm$ 3.46 98.7 $\pm$ 0.42
NACP 9.0 $\pm$ 0.46 93.0 $\pm$ 0.49 4.8 $\pm$ 0.13 93.0 $\pm$ 0.48 2.5 $\pm$ 0.09 93.0 $\pm$ 0.52
CP (oracle) 14.9 $\pm$ 0.60 90.0 $\pm$ 0.61 6.9 $\pm$ 0.19 90.0 $\pm$ 0.62 3.8 $\pm$ 0.13 90.02 $\pm$ 0.58
Noisy-CP 99.7 $\pm$ 3.67 99.7 $\pm$ 0.08 101.4 $\pm$ 3.58 99.5 $\pm$ 0.09 98.3 $\pm$ 3.80 99.8 $\pm$ 0.05
NR-CP (w/o $\Delta$)   14.0 $\pm$ 0.91   89.4 $\pm$ 0.81 6.7 $\pm$ 0.27   89.3 $\pm$ 0.80   3.5 $\pm$ 0.24   89.3 $\pm$ 0.80
TinyImagenet ACNL 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00
(200 classes) CRCP 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00 200.0 $\pm$ 0.00 100.0 $\pm$ 0.00
NACP 22.6 $\pm$ 1.87 93.7 $\pm$ 0.71 9.0 $\pm$ 0.50 93.6 $\pm$ 0.70 7.0 $\pm$ 0.87 93.6 $\pm$ 0.72
CP (oracle) 16.6 $\pm$ 0.33 90.0 $\pm$ 0.26 6.3 $\pm$ 0.06 90.0 $\pm$ 0.27 3.6 $\pm$ 0.07 90.0 $\pm$ 0.28
Noisy-CP 502.6 $\pm$ 8.56 99.9 $\pm$ 0.01 501.6 $\pm$ 8.51 99.9 $\pm$ 0.01 501.3 $\pm$ 10.2 100.0 $\pm$ 0.01
NR-CP (w/o $\Delta$) 16.7 $\pm$ 0.51   90.0 $\pm$ 0.34 6.3 $\pm$ 0.10   90.0 $\pm$ 0.36 3.6 $\pm$ 0.14   90.0 $\pm$ 0.38
ImageNet ACNL 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00
(1000 classes) CRCP 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00 1000.0 $\pm$ 0.00 100.0 $\pm$ 0.00
NACP 20.9 $\pm$ 0.72 91.9 $\pm$ 0.32 7.1 $\pm$ 0.13 91.9 $\pm$ 0.34 4.8 $\pm$ 0.23 91.9 $\pm$ 0.36
@@
$\textrm{size}=\frac{1}{n}\sum_{i}|C(x_{i})|,\hskip 8.5359pt\textrm{coverage}=\frac{1}{n}\sum_{i}{\textbf{1}}(y_{i}\in C(x_{i}))$
@@
Datasets. We show results on four standard scenery image datasets CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), Tiny-ImageNet, and ImageNet (Deng et al., 2009).
@@
Implementation details. Each task was trained by fine-tuning on a pre-trained ResNet-18 (He et al., 2016) network. The models were taken from the PyTorch site444 https://pytorch.org/vision/stable/models.html. We selected this network architecture because of its widespread use in classification problems. The last fully connected layer output size was modified to fit the corresponding number of classes for each dataset. For the standard dataset evaluated in Table 1 we used publicly available checkpoints. For each dataset, we combined the calibration and test sets and then constructed 1000 different splits where 50% was used for the calibration phase and 50% was used for testing. In all our experiments we used $\delta=0.001$. In other words, the computed coverage guarantee is applied to the sampled noisy calibration set with probability 0.999,